OpenStudy (anonymous):

Medal to the correct answer!!! One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? 4 min 5 min 6 min

OpenStudy (anonymous):

I think it's 5 but idk.

OpenStudy (agl202):

Ur close... A can do the job in 10 minutes or, In 1 minute, A can do 1/10 of the job. B can do the job in 15 minutes or, In 1 minute, B can do 1/15 of the job. Together, in 1 minute they can do 1/10+1/15 = 1/6 of the job. So, it would take 6/1 = 6 minutes to finish the job if they both work together.

OpenStudy (anonymous):

Thank you!

OpenStudy (agl202):

Ur welcome